OpenStudy (anonymous):

prove cos(3pi/2 )-A) =-sin A i am no good at these

OpenStudy (anonymous):

i wrote you a hint before :| why did you open a new one use the identity: cos(a+b) = cos(a)cos(b) - sin(a)sin(b) and another hint cos(3pi/2) = 0

OpenStudy (anonymous):

i dont understand what to do?

OpenStudy (anonymous):

cos(3pi/2-A) let 3pi/2 = x , -A = y now use cos(x+y) = cos(x)cos(y) - sin(x)sin(y) and see if you get -sin(A)

OpenStudy (phi):

cos(a+b) is short hand. You match the letter a to the first term in your problem and match the letter b to the 2nd term it might be easier if you use cos(a-b)= cos(a)cos(b)+sin(a)sin(b)

OpenStudy (anonymous):

got it thx

OpenStudy (anonymous):

took me a while didnt think to use my unit circle!